OpenStudy (anonymous):

convert the polar equation r/6=csc x to rectangular a

ganeshie8 (ganeshie8):

\(\large \frac{r}{6} = \csc \theta\)

ganeshie8 (ganeshie8):

like that ?

ganeshie8 (ganeshie8):

if so, il give a hint maybe : \(y = r \sin \theta\)

OpenStudy (anonymous):

Thanks !

ganeshie8 (ganeshie8):

np :) so wats teh rectangular equaiton u got ?